# Automated threshold selection and associated inference uncertainty for univariate extremes

### v3

@@ 4.2 Choice of tuning parameters
Values for the tuning parameters $(k,m)$ of the EQD and Varty methods are discussed and evidence is presented, based on an in-depth study in supplementary material S:5.3, to show that the methods are robust to their choice. Based on this evidence, unless stated otherwise, we take $(k,m)=(100,500)$ throughout the paper and in the supplementary material.
As $k$ controls the number of bootstrapped samples and thus, the level of sampling variability that is incorporated into the threshold choice, we expect higher values of $k$ to lead to more stable threshold choices. The RMSE values for threshold estimation reflect this, but also show that computation time increases linearly with $k$. For a one-off analysis, there is certainly merit in taking as large a value for $k$ that is computationally feasible. For simulation studies, when the computational implications of the choice of $k$ are more important, we find that $k=100$ balances accuracy and computation time sufficiently.
@@
4.3 EQD or Varty Method?
We compare the EQD and Varty methods and summarise the evidence of the extensive simulation study in supplementary material S:5.2. Specifically, we assess which case of metric (2) performs better for threshold selection and quantile estimation. In terms of threshold selection, we find that both methods perform similarly, with each method narrowly achieving the smallest RMSE in two of Cases 1-4, discussed in Section 6. However, for the estimation of high quantiles, the EQD outperforms the Varty method obtaining the lowest RMSE in the majority of cases and quantiles, due to having the smaller variance of estimates. Ultimately, we aim to estimate high quantiles accurately following a threshold selection, and so this study indicates that the EQD should be preferred, and so for the remainder of the paper, to avoid unnecessary repetition, we omit the results for the Varty method.
@@ 5 Accounting for parameter and threshold uncertainty
Even if the true threshold $u$ is known, relying on point estimates for the parameters of the fitted model results in misleading inference (Coles and Pericchi,, 2003). CIs are needed, but as standard errors and profile likelihoods rely on asymptotic arguments, they are not ideal due to the sparsity of threshold exceedances. We prefer parametric bootstrap methods which, as discussed in Section 2, give a similar performance to the profile likelihood for large samples. Algorithm 1 details the bootstrapping procedure to account for parameter uncertainty when $u$ is known. A GPD is fitted to the $n_{u}$ data excesses of $u$ from a sample $\bm{x}$ of size $n$ $(n\geq n_{u})$. Using the fitted parameters, $m_{1}$ parametric bootstrap samples above $u$ are simulated, each of size $n_{u}$, and the GPD parameters are re-estimated for each sample. A summary statistic, e.g., a return level, $s(u,\lambda_{u},\sigma_{u},\xi)$, may be computed for each of the $m_{1}$ bootstrapped values for $(\sigma_{u},\xi)$. This enables the construction of CIs for the GPD parameters and return levels [^1] Algorithm 1 focuses on the uncertainty of the estimates of $(\sigma_{u},\xi)$. Uncertainty in $\lambda_{u}$ estimates could be incorporated by simulating a number of excesses in each bootstrap sample from a Bin$(n,\lambda_{u})$ distribution. This uncertainty is included in Algorithm 2. .
@@
return A set of $m_{1}$ bootstrapped estimates for the summary statistic of interest.
GPD inferences are sensitive to the choice of threshold (Davison and Smith,, 1990) but uncertainty about this choice is not represented in Algorithm 1. This omission would be important when return levels inform the design of hazard protection mechanisms, where omitting this source of uncertainty could lead to overconfidence in the inference and have dangerous consequences. Algorithm 2 provides a novel method to propagate both threshold and parameter uncertainty through to return level estimation, using a double-bootstrap procedure. To focus on the threshold uncertainty and to forgo the need for a parametric model below the threshold, we employ a non-parametric bootstrap procedure on the original dataset. We resample with replacement $n$ values from the observed data $m_{2}$ times, estimate a threshold for each such bootstrap sample using the automated selection method of Section 4, and fit a GPD to the excesses of this threshold. For each one of the $m_{2}$ samples, we employ Algorithm 1 to account for the subsequent uncertainty in the GPD parameters. Calculating a summary statistic for each of the $m_{1}\times m_{2}$ samples leads to a distribution of bootstrapped estimates that accounts for both threshold and parameter uncertainty. We use $m_{1}=m_{2}=200$. In Section 6, we illustrate how using Algorithm 2 improves the coverage probability of CIs, and in Section 7 how it widens CIs for return levels of the River Nidd.
@@
     Employ Algorithm 1 with inputs: $(\bm{x}_{b},\hat{u}_{b},m_{1})$.
@@
return A set of $m_{1}\times m_{2}$ bootstrapped estimates for the summary statistic of interest.
@@ 6.1 Overview
We illustrate the performance of the EQD method against the Wadsworth, (2016) and Northrop et al., (2017) approaches, which we term the Wadsworth and Northrop methods respectively. Danielsson et al., (2001, 2019) approaches performed considerably worse than all others in threshold selection and quantile estimation; so results for these methods are only given in supplementary material S:6.2. We utilised the following R code for Wadsworth, Northrop and EQD methods respectively: Wadsworth, (2016) in the supplementary materials, threshr (Northrop and Attalides,, 2020), and https://github.com/conor-murphy4/automated_threshold_selection (Murphy et al.,, 2023).
We assess the methods’ ability to estimate the true threshold (when it exists) and the true return levels, using the root-mean-square error (RMSE) to measure performance. The true quantiles and all bias-variance components of RMSE, discussed in this section, are given in supplementary materials S:4 and S:6.2 respectively. We also investigate the merits of including the uncertainty in threshold selection in our inference, as discussed in Section 5, in terms of how they improve the coverage levels of CIs relative to their nominal values.
@@ 6.2 Scenario 1: True threshold for GPD tail
We consider Cases 0-8, with different properties above and below the true threshold of $u=1.0$ and various sample sizes. Case 0, where all of the data are from a GPD, is reported in supplementary material S:6.3, with the EQD performing notably better than the existing methods. Here, we present Cases 1-4, with Table 1 providing outline model and sample size properties, with full details and density plots given in supplementary material S:4. Cases 5-8 are presented in supplementary material S:6.3 with the EQD exhibiting the strongest performance and the Wadsworth method consistently failing for Cases 5-7; whereas Case 8, is Case 1 but with an unrealistic sample size of $n=20000$, shows that EQD performs comparably with the asymptotically justified Wadsworth method.
Cases 1-3 all have a distinct changepoint in the density at the true threshold, with Case 1, with $\xi>0$, being the simplest due to the clearly defined true threshold and large sample size. Case 2 is identical to Case 1 except for having a much smaller sample size. We find that the Wadsworth method failed to estimate a threshold in samples with $\xi<-0.05$ irrespective of sample size, so Case 3 is selected near that boundary where the method works reasonably, with $\xi=-0.05$ and double the sample size of Case 1. Case 4 provides a more difficult example with a continuous density and a small number of exceedances of the true threshold. The data are derived from a partially observed GPD, denoted GPDp, with data from a GPD above 0 and rejected if less than an independent realisation from a Beta(1,2) distribution.
For each case, the results are based on 500 replicated samples, for which we test the candidate thresholds $\{u_{i}\}_{i=1}^{20}$, corresponding to the sample quantile levels of $0\%,5\%,\ldots,95\%$, with the true threshold being the $16.67\%$ quantile for Cases 1-3 and the $72.10\%$ quantile for Case 4.
@@
Threshold recovery: Table 2 shows the RMSE of the chosen thresholds for each method in Cases 1-4. The EQD achieves RMSEs 1.2-7.7 (1-11.2) times smaller than the Wadsworth (Northrop) method. The EQD has the lowest bias in $3/4$ of cases and the least variance in all cases, improving in terms of bias (variance) by a factor between 0.8-6.6 (19-109) over the Wadsworth method and 0.4-8.1 (20-211) over the Northrop method.
@@
Quantile recovery: Table 3 presents the RMSEs for the $(1-p_{j})$-quantiles where $p_{j}=1/(10^{j}n)$ for $j=0,1,2$ for sample size $n$, which ensures that extrapolation is equally difficult over $n$ for a given $j$. When $j=0$, no extrapolation is required so the choice of $u$ should not be too important; the similar RMSEs across methods reflect this. As $j$ increases, all RMSEs increase and the differences between methods become clear. The EQD method is best uniformly, followed by the Wadsworth and then the Northrop methods. This pattern reflects the findings in Table 2, although here with differential performances sensitive to $j$. The EQD achieves the lowest bias in the majority of cases and quantiles with considerably less variance in quantile estimates in all cases, particularly as $j$ increases.
EQD Wadsworth2 Northrop EQD Wadsworth2 Northrop
@@
True quantile coverage: We apply Algorithms 1 and 2 to data from Case 4, the hardest case for threshold selection. Table 4 presents the coverage probabilities of the nominal 80% and 95% CIs of the estimated $(1-p_{j})$-quantiles as well as the average ratio of the CI widths over the 500 samples, termed CI ratio. Results for extra quantile levels, as well as coverage for the 50% CI, are given in supplementary material S:7. Overall, incorporating only parameter uncertainty (Alg 1) leads to underestimation of interval widths and inadequate coverage of the true quantiles, especially as we extrapolate further. The inclusion of the additional threshold uncertainty (Alg 2) leads to much more accurate coverage of the true quantiles across all exceedance probabilities. The CI ratios show that this highly desirable coverage is achieved with only 40-60% increase in the CI widths on average.
@@
CI ratio 1.420 1.443 1.467 1.461 1.524 1.602
Table 4: Coverage probabilities for estimated quantiles using Algorithm 1 and 2 for 500 replicated samples from Case 4 with sample size of 1000.
@@ 6.3 Scenario 2: Gaussian data
In applications there is no true or known value for the threshold above which excesses follow a GPD, so we explore this case here. We select the standard Gaussian distribution as it has very slow convergence towards an extreme value limit (Gomes,, 1994), so threshold selection is likely to be difficult. We assess threshold selection methods based on estimation of the true quantiles $\Phi^{-1}(1-p_{j})$ where $p_{j}=1/(10^{j}n)$, for $j=0,1,2$. We simulate 500 samples, for $n=2000$ and $20000$, with candidate thresholds being the $50-95\%$ sample quantiles, in steps of 5% and 0.5% for the two choices of $n$ respectively. As commented on for Case 8, in Section 6.2, $n=20000$ is unrealistic, but we include it to illustrate the slow convergence.
Quantile recovery: Table 5 shows the RMSEs of the estimated quantiles. For $n=2000$, the EQD method achieves the smallest RMSE with the Northrop method a close second, with the reverse when $n=20000$. The median and 95% CI of the chosen thresholds are given in supplementary material S:6.2. The Northrop method tends to choose slightly higher thresholds than the EQD method, leading to a small reduction in bias but for only the smaller $n$ is the additional variability relative to the EQD a disadvantage. The Wadsworth method performs worst as it incurs the most bias due to the lower threshold choices.
@@
Table 5: RMSEs of estimated $(1-p_{j})$-quantiles for 500 replicated samples from a Gaussian distribution for samples of size $n$. The smallest RMSE are highlighted in bold.
True quantile coverage: For assessing the coverage of true quantiles using Algorithms 1-2 for Gaussian data, Table 6 presents the coverage probabilities of the nominal 80% and 95% CIs of the estimated quantiles, when $n=2000$, as well as the average ratio of the CI widths over the 500 samples, with more results given in supplementary material S:7. Across the $p_{j}$, Alg 1 gives very low coverage probabilities in both cases, with performance deteriorating as $j$ increases. The added threshold uncertainty from Alg 2 results in large increases in coverage though still somewhat less than required, with this achieved through increases in CI widths by 40-70% on average. This weaker performance than we find in Section 6.2 suggests that no sample threshold (for realistic sample sizes) is large enough to overcome bias in making extreme value approximations for Gaussian data, but the improvement in coverage using Alg 2 demonstrates the importance of including the additional threshold uncertainty.
@@
CI ratio 1.444 1.469 1.498 1.464 1.548 1.639
Table 6: Coverage probabilities for estimated $(1-p)$-quantiles using Algorithm 1 and 2 for 500 replicated samples from a Gaussian distribution with sample size of 2000.

### v4

@@ 4.3 Choice of tuning parameters
An in-depth study in supplementary material S:5.3 demonstrates that the EQD method is robust to the choice of the tuning parameters $B$ and $m$. Consequently, we take $(B,m)=(100,500)$ throughout the paper and in the supplementary material, unless stated otherwise.
The number of bootstrapped samples $B$ controls the level of sampling variability that is incorporated into the threshold choice and so we expect higher values of $B$ to lead to more stable threshold choices. The RMSE values for threshold estimation reflect this but also show that computation time increases linearly with $B$. For a one-off analysis, there is certainly merit in taking as large a value for $B$ that is computationally feasible. For simulation studies, when the computational implications of the choice of $B$ are more important, we find that $B=100$ balances accuracy and computation time sufficiently.
The tuning parameter $m$ gives the number of equally-spaced evaluation probabilities used in expression (2). The EQD metric aims to approximate the IAE between model quantiles and quantiles of the data generating process (i.e., not for a particular sample) and a larger choice of $m$ improves this approximation. To compare fairly across a range of candidate thresholds, we choose to keep the quality of the approximation of the IAE fixed across thresholds and bootstraps by fixing the number $m$ of points in the quantile interpolation grid.
@@ 5 Accounting for parameter and threshold uncertainty
Even if the true threshold $u$ is known, relying on point estimates for the GPD parameters results in misleading inference (Coles and Pericchi,, 2003). CIs are needed, but as standard errors and profile likelihoods rely on asymptotic arguments, they are not ideal due to the sparsity of threshold exceedances. We prefer parametric bootstrap methods which, as discussed in Section 2, perform similarly to the profile likelihood for large samples. Algorithm 1 details the bootstrapping procedure to account for GPD parameter uncertainty when $u$ is known. A GPD is fitted to the $n_{u}$ data excesses of $u$ from a sample $\bm{x}$ of size $n$ $(n\geq n_{u})$. Using the fitted parameters, $B_{1}$ parametric bootstrap samples above $u$ are simulated, each of size $n_{u}$, and the GPD is re-estimated for each sample. A summary statistic, e.g., a return level, $s(u,\lambda_{u},\sigma_{u},\xi)$, may be computed for each of the $B_{1}$ bootstrapped values for $(\sigma_{u},\xi)$. This enables the construction of CIs for the GPD parameters and return levels.
@@
return A set of $B_{1}$ bootstrapped estimates for the summary statistic of interest.
Algorithm 1 focuses on the uncertainty of the estimates of $(\sigma_{u},\xi)$. We incorporate the additional uncertainty in the estimation of $\lambda_{u}$ by replacing the fixed $n_{u}$ in the loop over $b$ with a random variate from a Bin$(n,\hat{\lambda}_{u})$ distribution for each bootstrap sample, with this extension then referred to as Algorithm 1b.
GPD inferences are sensitive to the choice of threshold (Davison and Smith,, 1990) but uncertainty about this choice is not represented in Algorithms 1 or 1b. This omission would be important when return levels inform the design of hazard protection mechanisms, where omitting this source of uncertainty could lead to over-confidence in the inference and have dangerous consequences. Algorithm 2 provides a novel method to propagate both threshold and parameter uncertainty through to return level estimation, using a double-bootstrap procedure. To focus on the threshold uncertainty and to forgo the need for a parametric model below the threshold, we employ a non-parametric bootstrap procedure on the original dataset. We resample with replacement $n$ values from the observed data $B_{2}$ times, estimate a threshold for each such bootstrap sample using the automated selection method of Section 4, and fit a GPD to the excesses of this threshold. For each one of the $B_{2}$ samples, we employ Algorithm 1 to account for the subsequent uncertainty in the GPD parameters. Calculating a summary statistic for each of the $B_{1}\times B_{2}$ samples leads to a distribution of bootstrapped estimates that accounts for uncertainty in the threshold selection as well as in the GPD and threshold exceedance rate parameters. We use $B_{1}=B_{2}=200$. In Section 6, we illustrate how using Algorithm 2 improves the coverage probability of CIs, and in Section 7 how it widens CIs for return levels of the River Nidd.
@@
     Employ Algorithm 1 with inputs: $(\bm{x}_{b},\hat{u}_{b},B_{1})$.
@@
return A set of $B_{1}\times B_{2}$ bootstrapped estimates for the summary statistic of interest.
@@ 6.1 Overview
We illustrate the performance of the EQD method against the Wadsworth, (2016) and Northrop et al., (2017) approaches, which we term the Wadsworth and Northrop methods respectively. Danielsson et al., (2001, 2019) approaches perform considerably worse than all others in threshold selection and quantile estimation; so results for these methods are only given in supplementary material S:6.2. We utilised the following R code for Wadsworth, Northrop and EQD methods respectively: code given in the supplementary materials of Wadsworth, (2016), threshr (Northrop and Attalides,, 2020), and https://github.com/conor-murphy4/automated_threshold_selection (Murphy et al.,, 2023).
The performance of all of the methods depends somewhat on the choice of the set of candidate thresholds which we denote by:
$C_{u}=\{u_{i},i=1,\ldots,k:u_{1}<\ldots<u_{k}\},$ (3)
where we restrict the $u_{i}$ to be sample quantiles evaluated at equally-spaced probabilities. The range $[u_{1},u_{k}]$, the number of candidates $k$ and the inter-threshold spacing are all potentially important in terms of how they affect the performance of the methods. As emphasised in Section 1, we are aiming for an automated threshold selection method which can achieve accurate results without any user inputs, so a key element of our study is to investigate how these features of the set $C_{u}$ impact on the methods’ relative performance. When fitting a GPD with decreasing density (i.e., for $\xi>-1$), it would be inadvisable to use a threshold which clearly lies below the mode of the distribution. As we want to avoid the requirement of user estimation of the mode, our standard choice for the range of the candidate grid is $[u_{1},u_{k}]$: $(u_{1},u_{k})=(0\%,95\%)$ sample quantiles of all the data. However, we also explore several cases where only the data lying above the mode are used with $[u_{1},u_{k}]$: $(u_{1},u_{k})=(0\%,95\%)$ now sample quantiles of the remaining data. To remove the uncertainty arising from the choice of estimator of the mode, we use the true mode which has a unique value in our simulated cases. Results in supplementary material S:6.4 indicate that our original choice for the candidate threshold set does not unfairly favour the EQD method in any way.
We consider two scenarios: Scenario 1 and Scenario 2 where the true threshold is known and unknown respectively. Here, we present the results using a candidate threshold grid across the whole distribution for Scenario 1 and above the sample median for Scenario 2, with the latter chosen as the Wadsworth method fails when applied across the default range in that setting. The Wadsworth method relies on asymptotic arguments, which limits how large $k$ can be relative to the sample size, $n^{\prime}$, above the mode, with $n^{\prime}\leq n$, where $n$ is the total sample size. To assess how the Wadsworth method performs as a fully automated method, we apply the method despite the value of $k$ not always aligning with the guidance in Wadsworth, (2016) about its size relative to $n^{\prime}$.
We assess the methods’ ability to estimate the true threshold (when it exists) and the true return levels, using the RMSE to measure performance. The true quantiles and all bias-variance components of RMSE, discussed in this section, are given in supplementary materials S:4 and S:6.2 respectively. We also investigate the merits of including the uncertainty in threshold selection in our inference, as discussed in Section 5, in terms of how they improve the coverage levels of CIs relative to their nominal values.
@@ 6.2 Scenario 1: True threshold for GPD tail
We consider Cases 0-8, with different properties above and below the true threshold of $u=1.0$ and various sample sizes. Case 0, where all of the data are from a GPD, is reported in supplementary material S:6.3, with the EQD performing notably better than the existing methods. Here, we present detailed results for Cases 1-4, with Table 1 providing outline model and sample size properties, with full details and density plots given in supplementary material S:4. Cases 5-8 are considered briefly after discussing Cases 1-4 below.
Cases 1-3 all have a distinct changepoint in the density at the true threshold. Cases 1 and 2 have the same distribution, with $\xi>0$, with Case 2 having a smaller sample size. We find that the Wadsworth method fails to estimate a threshold in samples with $\xi<-0.05$ irrespective of sample size, so Case 3 is selected near that boundary where the method works reasonably and has double the sample size of Case 1. Case 4 provides a more difficult example with a continuous density and a small number of exceedances of the true threshold. The data are derived from a partially observed GPD, denoted GPDp, with data from a GPD above 0 and rejected if less than an independent realisation from a Beta(1,2) distribution.
For each case, the results are based on 500 replicated samples, for which we test the candidate thresholds $C_{u}$, with $k=20$, as given in (3), with the true threshold being the $16.67\%$ quantile for Cases 1-3 and the $72.10\%$ quantile for Case 4.
@@
Cases 1-4, Threshold recovery: Table 2 shows the RMSE of the chosen thresholds for each method in Cases 1-4. The EQD achieves RMSEs 1.2-7.7 (1-11.2) times smaller than the Wadsworth (Northrop) method. The EQD has the lowest bias by a considerable margin in Cases 1-3 and shows the lowest variance in threshold estimation in all cases. In fact, the variance is reduced by factor of at least 20 relative to both the Wadsworth and Northrop methods (see Table S.9).
@@
Cases 1-4, Quantile recovery: Table 3 presents the RMSEs for the $(1-p_{j,n})$-quantiles where $p_{j,n}=1/(10^{j}n)$ for $j=0,1,2$ for sample size $n$, which ensures that extrapolation is equally difficult over $n$ for a given $j$. When $j=0$, no extrapolation is required so the choice of $u$ should not be too important; the similar RMSEs across methods reflect this. As $j$ increases, all RMSEs increase and the differences between methods become clear. The EQD method is best uniformly, followed by the Wadsworth and then the Northrop methods. This pattern reflects the findings in Table 2, although here with differential performances sensitive to $j$. The EQD achieves the lowest bias in the majority of cases and leads to quantile estimates with considerably less variance in all cases, particularly as $j$ increases.
EQD Wadsworth1 Northrop EQD Wadsworth1 Northrop
@@
Summary for Cases 5-8: Cases 5-8 are very similar in form to Cases 1-3 but with different shape parameters and sample sizes. The results for these cases are presented in supplementary material S:6.3, with a brief summary given here. Specifically, for Cases 5-7, we find that the EQD exhibits the strongest performance and the Wadsworth method consistently fails due to the small sample sizes or computational issues with numerical integration when $\xi<-0.05$. Case 8 is parameterised similarly to Case 1 but with an unrealistic sample size of $n=20000$. Although the data in Case 8 are more suited to a method reliant on asymptotic theory, the EQD performs comparably with the Wadsworth method, with both performing better than the Northrop method.
Case 4, True quantile coverage: We apply Algorithms 1, 1b and 2 to data from Case 4, the hardest case for threshold selection. Table 4 presents the coverage probabilities of the nominal 80% and 95% CIs of the estimated $(1-p_{j,n})$-quantiles as well as the average ratio of the CI widths (based on Alg 2 relative to Alg 1) over the 500 samples, termed CI ratio. Results for extra quantile levels, as well as coverage for the 50% CI, are given in supplementary material S:7. Overall, incorporating only parameter uncertainty (Alg 1 and Alg 1b) leads to underestimation of interval widths and inadequate coverage of the true quantiles, especially as we extrapolate further. The additional uncertainty, given in Alg 1b, by also accounting for uncertainty in the rate of threshold exceedance, typically makes a very small improvement in coverage, and for some quantiles, this actually leads to a reduction in coverage due to Monte Carlo variation in the simulations. In contrast, the inclusion of the additional threshold uncertainty (Alg 2) leads to much more accurate coverage of the true quantiles across all exceedance probabilities. The CI ratios show that this highly desirable coverage is achieved with only 43-62% increase in the CI widths on average.
@@
CI ratio 1.430 1.452 1.475 1.484 1.546 1.621
Table 4: Coverage probabilities for estimated quantiles using Algorithms 1, 1b and 2 for 500 replicated samples from Case 4 with sample size of 1000. CI ratio gives the average ratio of the CIs for Algorithm 2 relative to Algorithm 1 over the 500 samples.
@@ 6.3 Scenario 2: Gaussian data
In applications, there is no true or known value for the threshold above which excesses follow a GPD, so we explore this case here. We select the standard Gaussian distribution as it has very slow convergence towards an extreme value limit (Gomes,, 1994), so threshold selection is likely to be difficult. We assess threshold selection methods based on estimation of the true quantiles $\Phi^{-1}(1-p_{j,n})$ where $p_{j,n}=1/(10^{j}n)$, for $j=0,1,2$. We simulate 500 samples, for $n=2000$ and $20000$, with $C_{u}$, given in (3), now having range $[u_{1},u_{k}]:(u_{1},u_{k})=(50\%,95\%)$ sample quantiles of the data and $k=10$ and 91 (i.e., steps of 5% and 0.5%) for the two choices of $n$ respectively. As with Case 8 in Section 6.2, $n=20000$ is unrealistic, but we include it to illustrate the slow convergence.
Quantile recovery: Table 5 shows the RMSEs of the estimated quantiles. For $n=2000$, the EQD method achieves the smallest RMSE with the Northrop method a close second, with the reverse when $n=20000$. The median and 95% CI of the chosen thresholds are given in supplementary material S:6.2. The Northrop method tends to choose slightly higher thresholds than the EQD method, leading to a small reduction in bias, but for only the smaller $n$ is the additional variability relative to the EQD a disadvantage. The Wadsworth method performs the worst, selecting lower thresholds and so incurring the most bias.
@@
Table 5: RMSEs of estimated $(1-p_{j,n})$-quantiles for 500 replicated samples from a Gaussian distribution for samples of size $n$. The smallest RMSE are highlighted in bold.
True quantile coverage: For assessing the coverage of true quantiles using Algorithms 1, 1b and 2 for Gaussian data, Table 6 presents the coverage probabilities of the nominal 80% and 95% CIs of the estimated quantiles, when $n=2000$, as well as the average ratio of the CI widths (again, of Alg 2 relative to Alg 1) over the 500 samples, with more results given in supplementary material S:7. Across the $p_{j,n}$, both Alg 1 and 1b give very low coverage probabilities in both cases, with performance deteriorating as $j$ increases. The added threshold uncertainty from Alg 2 results in large increases in coverage though still somewhat less than required, with this achieved through increases in CI widths by 45-66% on average. This weaker performance than we find in Section 6.2 suggests that no sample threshold (for realistic sample sizes) is large enough to overcome bias in making extreme value approximations for Gaussian data, but the improvement in coverage using Alg 2 demonstrates the importance of including the additional threshold uncertainty.
@@
CI ratio 1.457 1.480 1.509 1.495 1.576 1.665
Table 6: Coverage probabilities for estimated quantiles using Algorithms 1, 1b and 2 for 500 replicated samples from a Gaussian distribution with sample size of 2000. CI ratio gives the average ratio of the CIs for Algorithm 2 relative to Algorithm 1 over the 500 samples.